When you measured your diagonal clothesline across your rectangular porch, you found that it was 1 foot longer than the length of the porch, and 8 ft longer than the width. What are the dimensions of your porch?
Thanks for the help!!

Answer by checkley77(12844)

(x-1)^2+(x-8)^2=x^2
x^2-2x+1+x^2-16x+64=x^2
x^2-18x+65=0
(x-13)(x-5)=0
x-13=0
x=13 ans.
Proof:
(13-1)^2+(13-8)^2=13^2
12^2+5^2=169
144+25=169
169=169